Four capacitors are connected as shown below. Fin
Answer #1

(a) Equivalent capacitance is calculated as follows

15 uF and 3 uF are in series ,their net=45/18=5/2=2.5 uF ,and this is now in paralle with 6 uF so net =8.5 uF

now this 8.5uF is in series with 20 uF

so net capacitane=8.5*20/(8.5+20)=5.97 uF

(b)net charge=CV =5.97*15=89.47 uC

charge on 20 uF=89.47 uC

charge on 6uF=(6/8.5)*89.47=63.16 uF

charge on 15 uF=charge on 3 uF=(89.47-63.16)=26.31 uF

(c) energy of 3 uF=0.5Q^2/3uF=1.15*10^(-4) J